A mercury thermometer is constructed as
shown. The capillary tube has a diameter
of 0.0037 cm, and the bulb has a diameter of
0.28 cm. Neglecting the expansion of the glass, find
the change in height of the mercury column
for a temperature change of 41 ◦C. The
volume expansion coefficient for mercury is
0.000182 (◦C)−1.Answer in units of cm.
1
Expert's answer
2020-03-10T11:17:46-0400

The volume change of the mercury is


"\u2206V=\u03b2 V\u2206T"

and it should be equal to the mercury volumein the tube, since we ignore the expansion of the glass


"\u2206h A= \u2206V=\u03b2 V\u2206T"

"\pi r^2∆h=\frac{4\pi R^3}{3}β ∆T"

"(0.00185)^2∆h=\frac{4 (0.14)^3}{3}(0.000182)(41)"

"∆h=7.98\ cm"